OpenStudy (96mertcank96):

the product of x^0, 4x^-2, and 5x^3

OpenStudy (anonymous):

Anything raised to power 0 except 0 is 1.. So tell me what is \(x^0\) = ??

OpenStudy (96mertcank96):

1

OpenStudy (anonymous):

Yep...

OpenStudy (anonymous):

\[x^{-a} = \frac{1}{x^a}\]

OpenStudy (96mertcank96):

i got up to 20x * x^0

OpenStudy (anonymous):

x^0 = 1 so what are you left with??

OpenStudy (96mertcank96):

yeah but there is no answer choice with 21x

OpenStudy (anonymous):

Hey, \(20x \times 1 = ??\)

OpenStudy (96mertcank96):

nice... thank you!

OpenStudy (anonymous):

You are multiplying 20x with 1 and not adding 1, here you can't add 1 too..